OpenStudy (anonymous):

Which value is NOT a solution of 8x3 – 1 = 0?

OpenStudy (anonymous):

what are the choices?

OpenStudy (anonymous):

A)1-i√3/4 B)2 C)1+i√3/4 D)1/2

OpenStudy (16shuston):

solve for x

OpenStudy (anonymous):

You sure that the question is right?

OpenStudy (anonymous):

yes thats how the question was written

OpenStudy (anonymous):

I only get D as a solution ._.

OpenStudy (16shuston):

ya same

OpenStudy (anonymous):

thats what i got but i wasn't sure thanks for the help everyone

jimthompson5910 (jim_thompson5910):

Let's test choice D 8x^3 - 1 = 0 8(1/2)^3 - 1 = 0 ... replace x with 1/2 8*(1/8) - 1 = 0 1 - 1 = 0 0 = 0 ... this is true so x = 1/2 is indeed a true solution to 8x^3 - 1 = 0

jimthompson5910 (jim_thompson5910):

They are asking for non-solutions.